Sol Summary – April 5th

Report title: Sol Summary Report
Crew #: 333
Position: Crew Commander
Report prepared by: Marie Jansen
Date: 05-04-2026
Sol: SOL 7

Summary Title: Easter Day
Mission Status: Breaking the routine on Easter Day.
Sol Activity Summary:
Today was our rest day. We decided to celebrate Easter by sharing each other’s traditions. We woke up at 9:00 AM and had chocolate chip pancakes for breakfast. After that we had an egg hunt in the HAB. We caught the animal that had been eating our plants in the GreenHab. For the rest of the morning, Antoine and Marie did 45 minutes of cardio while the girls washed their hair and Matthias worked on his assignment. As for Matias, he was our cheerleader for the workout. For lunch we had couscous cooked by Zahraa which was delicious. In the afternoon we took a nap to rest properly and then got ready for crew photos. After a 1 hour photoshoot, we played cards while enjoying an Easter cake prepared by Joanna and Marie. We are currently working on our reports while Matthias is preparing dinner. Mac and cheese is on the menu.

Look Ahead Plan:
Tomorrow’s plan: we will be carrying out Joanna’s experiment at Cowboy Corner. The goal is to have everyone go through it on the same day. She has therefore decided to stay fairly close to the base. This will take up the whole day. Some of us have planned to do sport. We will also be running Zahraa’s experiment in the lower hab.

Anomalies in work:
None
Weather:
No EVA was carried out today, but the computers reported good conditions outside plenty of sunshine and a fairly pleasant temperature.
Crew Physical Status:
Everyone in the crew is doing well, the rest of the day did us a lot of good.
EVA:
No AVE today.
Reports to be filed:
Sol Summary
Operations Report
Greenhab Report
Journalist Report
Astronomy Report
EVA Requests
Daily photos
Mid-mission research report

Support Requested: none

Crew Photos April 5th>

Report title: Photos of the Day
Crew #: 333
Position: Crew Journalist
Report prepared by: Joanna Galloway
Date: 05-04-2026
Sol: 7

GreenHab Report – April 5th

Report title: GreenHab Report
Crew #: 333
Position: GreenHab Officer
Report prepared by: Matthias De Groote
Date: 05-04-2026
Sol: 7

Environmental control (fan & heater): Yes, Fan max : 80, Fan min : 75 ; Heater : 70 F°
Average temperatures (last 24h): 85.2 F°
Maximum temperature (last 24h): 101.6 F°
Minimum temperature (last 24h): 68.6 F°
Hours of supplemental light: 4:00
Daily water usage for crops: 2.4 gal
Daily water usage for research and/or other purposes: 0.26 gal
Water in the Tank (160 gal useful capacity): 142 gallons remaining
Time(s) of watering for crops: 10:00 am
Changes to crops: Beans are starting to flower.
Harvest: 21 g of arugula and 116 g of kale.
Support/supplies needed: No
Notes: The space rodent invader has been caught, we stay on our guards for future potential invaders.

Astronomy Report – April 5th

[category 

astronomy-report]

Report title: Astronomy Report
Crew #: 333
Position: Crew Astronomer
Report prepared by: Matias Ballivian Ascarrunz
Date: 05-04-2026
Sol: 7

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested: RCOS-16 & MDRS-WF
Objects to be Imaged this Evening: M64, M82
Images submitted with this report: None
Problems Encountered: None
MUSK OBSERVATORY
Solar Features Observed: N/A
Images submitted with this report: N/A
Problems Encountered: N/A

Astronomy Report – April 5th

[category 

astronomy-report]

Report title: Astronomy Report
Crew #: 333
Position: Crew Astronomer
Report prepared by: Matias Ballivian Ascarrunz
Date: 06-04-2026
Sol: 7

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested: RCOS-16 & MDRS-WF
Objects to be Imaged this Evening: M64, M82
Images submitted with this report: None
Problems Encountered: None
MUSK OBSERVATORY
Solar Features Observed: N/A
Images submitted with this report: N/A
Problems Encountered: N/A

Operations Report – April 5th

Report title: Operations Report
Crew #: 333
Position: Executive Officer / Crew Engineer
Report prepared by: Antoine Dubois
Date: 05-04-2026
Sol: 7

Non-nominal systems: Pressurised tunnels
Notes on non-nominal systems: We repaired the pressurised tunnel using 6/7 zip ties.

ROVERS
Spirit rover used: Not used
Hours: /
Beginning charge: /
Ending charge: /
Currently charging: No
Opportunity rover used: Not used
Hours: /
Beginning charge: /
Ending charge: /
Currently charging: Yes
Curiosity rover used: Not used
Hours: /
Beginning charge: /
Ending charge: /
Currently charging: No
Perseverance rover used: Not used
Hours: /
Beginning charge: /
Ending charge: /
Currently charging: No
General notes on rovers: None

Summary of Hab operations:
Water Use: 13.03 gallons
Main tank level: 247 gallons
Main water tank pipe heater (ON or OFF): OFF
Main water tank heater (ON or OFF): OFF
Toilet tank emptied (NO or YES): NO

Summary of internet: Connection used for research, listening to music, sending reports.
Summary of suits and radios: None were used
Summary of GreenHab operations:
Water use: 2.6 gallons
Heater (ON or OFF): ON, 70 F°
Supplemental light (hours of operation): 7/30 pm to 11/30 pm
Harvest (name, weight in grams): 21 g of arugula and 116 g of kale
Summary of ScienceDome operations: Today, we just had a photo shoot with the whole team.
Dual split (Heat or AC, On or Off or Automatic): AC, Automatic
Summary of RAM operations: We used 6/7 zip ties to repair the pressurized tunnels.
Summary of any observatory issues: None
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support:
We just put some trash bags in the RAMM.

Journalist Report – April 5th

Report title: Journalist Report

Crew #: 333

Position: Crew Journalist

Report prepared by: Joanna Galloway

Date: 05-04-2026

Sol: 7

Journalist Report Title (if applicable):

Today was a rest day for the Arsia Crew, after a week of doing EVA’s, it was a nice change of pace. We enjoyed a slow morning, eating pancakes and chatting. After checking on the greenhab, we discovered that we managed to trap the space mouse which had destroyed our crops. It so happened to be Easter today so Joanna had organised for the Crew an easter egg hunt in the Hab. Matias found the most eggs but was nice enough to share it with the rest of the team. The girls finally had time to try and wash their hair while still trying to limit their water use, it was a fun sight to see while some other team members of the Crew did some Cardio on the static bike. For lunch we had couscous which we ate with pleasure. Digestion knocked us out so we all went for a nap, we’re also still adapting to the time difference between earth and mars so it was good to catch up on some lost sleep. We still had plenty of time in the afternoon to make cupcakes and play a little uno. We also took the opportunity to do some crew pictures in the afternoon which turned out perfectly. Not much to report in the station today, we have installed a nice routine and are happy in our little cocoon. A day to breathe has been nice and we’re excited to explore more of Mars tomorrow with our EVA’s. Tonight’s plan is to eat a pasta bake and watch a movie to end our relaxed Sunday on a high note.

Operations Report – April 4th

Report title: Operations Report
Crew #: 333
Position: Executive Officer / Crew Engineer
Report prepared by: Antoine Dubois
Date: 04-04-2026
Sol: 6

Non-nominal systems: No issues to report

Notes on non-nominal systems: /

ROVERS

Spirit rover used: No

Hours: /
Beginning charge: /
Ending charge: /
Currently charging: No

Opportunity rover used: EVA #9

Hours: 266.7
Beginning charge: 100%
Ending charge: <10%
Currently charging: Yes

Curiosity rover used: EVA #9

Hours: 354.7
Beginning charge: 100%
Ending charge: 31%
Currently charging: Yes

Perseverance rover used: No

Hours: /
Beginning charge: /
Ending charge: /
Currently charging: No

General notes on rovers: None

Summary of Hab operations:

Water Use (see notes): 16.98 gallons
Main tank level (remaining gallons, see notes): 261 gallons
Main water tank pipe heater (ON or OFF): OFF
Main water tank heater (ON or OFF): OFF
Toilet tank emptied (NO or YES): NO

Summary of internet: Connection used for research, listening to music, and sending reports.

Summary of suits and radios:

  • Marie, Commander: Suits and radio 1

  • Antoine, Executive officer: Suits and radio 2

  • Matias Ballivian, Astronomer: Suits and radio 4

  • Joanna, Journalist: Suits and radio 6

Summary of GreenHab operations:

Water use: 2.75 gallons
Heater (ON or OFF): ON, 70 F°
Supplemental light (hours of operation): 7:30 pm to 11:30 pm
Harvest (name, weight in grams): N/A

Summary of ScienceDome operations: None

Dual split (Heat or AC, On or Off or Automatic): AC, automatic

Summary of RAM operations: The astronomer used wire cutters, cable and a soldering iron for his experiment.

Summary of any observatory issues: None

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: /

Astronomy Report – April 4th

[category  astronomy-report]

Report title: Astronomy Report
Crew #: 333
Position: Crew Astronomer
Report prepared by: Matias Ballivian Ascarrunz
Date:04-04-2026
Sol: 6

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested: MDRS-WF & RCOS-16
Objects to be Imaged this Evening: M86, NGC 7000, M104
Images submitted with this report: M51 (Whirlpool Galaxy). The image was processed by stacking exposures taken with the same filter and then combining the layers using RGB + Lum.
Problems Encountered: None

MUSK OBSERVATORY
Solar Features Observed: N/A
Images submitted with this report: N/A
Problems Encountered: N/A

Questions: For image processing, what techniques would you recommend to achieve sharper images with brighter, more balanced colors? I find that the submitted image appears overly blue.

EVA Report – April 4th

Report title: EVA Report
Crew #: 333
Position: Commander
Report prepared by: Marie Jansen
Date: 04-04-2026
Sol: 6

EVA #: 10

Purpose of EVA:
For Matias’ experiment we will go to the North ridge repeater to see the difference in reception of the different channels (#1 and #2). Then we will go to the Kissing Camel Ridge West, to compare the difference in reception compared to the North ridge, because they are at a similar distance from the HAB.

Start time: 2:00 PM
End time: 04:20 PM

Narrative: They started decompression at 1:55 PM, They headed out on foot towards North Ridge. They arrived at 2:35 PM. They explored the area and found ways to communicate with us on both channels. They left North Ridge at 3:07 PM. They then headed towards Kissing Camel Ridge West (KCRW), but they confused the ridge with another so they arrived 500m north from KCRW. They departed from there at 4:07 PM. and arrived at the station back at 4:25 PM.

Destination per MDRS Map: North Ridge, Kissing Camel Ridge West

Destination Coordinates (use UTM WGS 84): 0518491E 4251750N; 0517819E 4249635N

EVA Participants:
Matias, Astronomer (Team lead)
Antoine, Executive Officer

Road(s) and routes per MDRS Map: Road 1103 and Hab Ridge Road

Mode of travel: Walking